18 April 2014 - Mission Day: 6713 - DOY: 108
Recovery
 
 

BLOGS: Recovery Updates (13 July 1998 - 4 November 1998)

4 November 1998

    The spacecraft status is nominal.

    SUMER successfully completed its recommissioning. All seven mechanisms are working perfectly as before. The detector gain is nominal. Detector B has been fully recalibrated. No sensitivity loss could be measured. The sensitivity is exactly as before. Recalibration of detector A is ongoing. The rear slit camera is working, too.

    This completes the recommissioning phase of the SOHO instruments. All twelve instruments are back to normal and fully functional (except the LASCO C1 coronagraph which is still being tested). Again my compliments to the scientists and engineers who designed and built these instruments and this spacecraft. They can really be proud of their workmanship.

    I think this is also a good time to stop with these daily reports, in particular since we have resumed circulating minutes of our daily planning meetings which are also available on the web under the SOHO Operations pages.

    It's still a miracle for me ...

3 November 1998

    The spacecraft status is nominal.

    SUMER check out of its mechanisms is continuing. So far 3 out of the 7 mechanisms have been checked and are working fine.

    The Spartan 201 satellite was successfully captured by the Shuttle Discovery at 20:48 UT today. All indications are that the mission proceeded flawlessly and is a complete success.

2 November 1998

30 October 1998

    The spacecraft status is nominal.

    CELIAS and MDI have finished their recommissioning. Both have been declared nominal and fully operational.

    The FOT has no remaining payload TSTOLs to upload. The next instrument re-certification activities are scheduled for November 2nd (SUMER).

29 October 1998

    The spacecraft status is nominal.

    MDI finished uploading the detector flat field.

    CDS has done a full Sun imaging test to calibrate the pointing of the instrument.

    CELIAS has all the sensors in nominal mode and fully operational. SEM was alredy declared re-certified earlier this week. PM, MTOF and STOF science data are being evaluated. CTOF is as good as it was before the loss of attitude.

28 October 1998

    The spacecraft status is nominal.

    CELIAS re-certification continues without problems. CELIAS/MTOF has been switched on today, so now all CELIAS sensors are on. As this was the only sensor on the spacecraft that remained off, all the detectors on board SOHO are powered on again.

27 October 1998

    The spacecraft status is nominal.

    CELIAS recovery is proceeding as scheduled. CELIAS/CTOF housekeeping parameters are nominal and the high voltage settings have been increased once more.

    CELIAS/STOF/HSTOF also shows nominal housekeeping, the solid state detectors and the HSTOF deflection system are now at operational levels. The microchannel plate bias has been enabled and set at check-out voltage levels.

    CELIAS/SEM science data appear to be of excellent quality and show that no degradation of the SEM has occurred during the last few months. From Toni Galvin: "The measured solar flux ratio between the central order channel and the two side channels is about what it was, for quiet sun conditions, before the temporary loss of spacecraft data, and the absolute flux values are currently somewhat higher than they were at the end of June, which is quite reasonable due to the fact that we are on the rising part of the solar activity cycle. The measured flux increase during the limited data received thus far is not great enough, however, to know whether it corresponds to short term variations or a real increase associated with the eleven year solar cycle."

26 October 1998

    The spacecraft status is nominal.

    CELIAS/STOF/HSTOF was powered on into manual mode on Sunday. All housekeeping parameters are nominal.

    CELIAS/SEM was also powered on yesterday and is fully operational. Housekeeping is nominal, and after the evaluation of the science data completed today it has been declared recommissioned.

    CELIAS/CTOF was switched on and placed into manual mode on Sunday, too. A calibration sequence was sucessfully completed. All high voltage supplies have been tested and initial voltages set. The microchannel plates are still somewhat cold, but their temperature is increasing gradually as the nearby electronics warm them up.

    The recovery status of CELIAS can be followed at http://umtof.umd.edu/~galvin/celias_recovery/.

24 October 1998

    The spacecraft status is nominal.

    CELIAS Digital Processing Unit (DPU) was successfully turned on at 17:56 UT. DPU software patches were successfully loaded. Tomorrow the CELIAS STOF sensor unit turn on is planned.

23 October 1998

    The spacecraft status is nominal.

    As of today, 23 October 1998, after nearly 4 months since the loss of contact with SOHO on 25 June, we have 10 of the 12 instruments working nominally again (or close to nominal nearing the end of their recommissioning phase): GOLF, VIRGO, MDI, CDS, EIT, LASCO (except C1 which is still being tested), UVCS, SWAN, COSTEP, ERNE. The last instrument to be switched back on is CELIAS. It will have its day tomorrow, 24 October. SUMER will resume recommissioning on 2 November.

22 October 1998

    The spacecraft status is nominal.

    The Swan instrument is now fully recommissioned and will resume its nominal observing program on Oct 23. The -Z side sensor is fully nominal. The photometric count rate and Hydrogen absorption levels for this sensor have been found to be similar to the values they had in June 1998. The +Z sensor shows a change in the counting efficiency. The detector seems to be more sensitive than before at larger wavelengths. This means that the calibration procedures and curves for this unit will have to be reestablished. The first skymap measured yesterday with this sensor shows all the expected features. So all the previous scientific objectives in photometric mode should be maintained for both sensors. Hydrogen cell scientific measurements seem also to be recoverable, after a serious calibration effort.

    The LASCO C1 camera mechanisms were tested and found to be working properly. Closed door C1 images show blur indicating that the Fabry-Perot filter plates are slightly misaligned. Continued testing of the C1 Fabry-Perot filter revealed response from all three piezos. Calibration data have been acquired to determine the wedge defection direction and magnitude.

    The VIRGO team had a successful balloon flight in Southern France with measurements of the total and spectral solar irradiance during 45 minutes from an altitude of about 42 km.

21 October 1998

    The spacecraft status is nominal.

    CDS opened its doors at 20:05 UT and acquired first light images both from the Normal Incidence Spectrometer (NIS) and Grazing Incidence Spectrometer (GIS). Both spectra look completely nominal. They are available at the CDS recommissioning page at URL: http://orpheus.nascom.nasa.gov/cds/diary/status.html

    MDI tested their image stabilization system (ISS). UVCS happened to be rolling at the time, providing a convenient jitter source with which to test the system -- both open and closed loop. Several amplifier gains were tested; the nominal gain still appears to work the best. The ISS was left in open loop (safe) mode overnight but is in good working order.

20 October 1998

    The spacecraft status is nominal.

    The CDS NIS detector is working.
    Several dark exposures were made with the microchannel plate (MCP) switched off, these showed the normal particle events. Five more exposures were then made with the MCP ON and the hot spot, that we have seen since before launch, was visible, confirming that the NIS detector is working. For details see http://orpheus.nascom.nasa.gov/cds/dairy/status.html

    SWAN:
    As of today 20:00 UT, the -Z sensor is fully recommissionned. The count rate and Hydrogen absorption levels for this sensor have been found to be similar to the values they had in June 1998. Normal operations can resume for this sensor.

    An anormal behaviour was found for the +Z sensor when the hydrogen cell was activated. The count rate increased for 24 out of the 25 pixels for a yet unknown reason. The 25th pixel shows an absorption pattern similar to what it was in June 1998. Tests of the +Z sensor hydrogen cell and detector are continuing.

    MDI:
    The MDI recommissioning is proceeding smoothly. Last night we turned on the MDI filter oven and adjusted the temperature to the standard operating point of 35 deg C. During the warm up of the filter oven, the filter tuning was changing and the dopplergrams shifted in intesting patterns - see for example http://mdiwww.nascom.nasa.gov/health_mon/gif/loi_img_hours.gif.

    After the oven temperature stabilized, the MDI dopplergrams, continuum images and magnetograms appear very similar to those taken earlier this year. The Michelson tuning parameters used in the reload appear to be close to optimal. This indicates that the filter performance does not appear to have substantially changed due to the extended cold-soak. Evaluation of calibration images taken during the oven warm-up will determine if further adjustment of the tuning is needed.

    Sample images taken with the oven at operating temperature can be found at:
    Dopplergram: http://mdiwww.nascom.nasa.gov/gif_recovery/fd_V.gif
    Continuum: http://mdiwww.nascom.nasa.gov/gif_recovery/fd_Ic.gif
    Magnetogram: http://mdiwww.nascom.nasa.gov/gif_recovery/fd_M.gif

    At this time, there are no known problems with the MDI instrument, and the MDI team should be able to return to standard observing shortly.

19 October 1998

    The spacecraft status is nominal.

    CDS: The four mechanisms were excercised again today and are nominal. The instrument temperatures are being slowly lowered to our nominal operating temperatures.(Duty cycles 60%) The NIS detector electronics subsystem was switched on and succesfully tested. The GIS detector subsystem was switched on and succesfully tested. The HV's to the four detectors was successfully ramped up to 3KV. Further HV tests are continuing.

    Test of all four SWAN mechanisms were sucessfully completed today. All mechanisms behave nominally. High voltage of the +Z SWAN sensor was switched on at 20:30 UT and slowly ramped up to its usual setting used before the mission interruption. A first analysis of the count rate shows a decrease of roughly a factor of 2 from what it was before. This can be compensated by increasing high voltage setting. High voltage of the -Z sensor was switched on at 21:00 UT. The count rate of this sensor is identical to what it was prior to mission interruption. Testing of both Hydrogen cells will be performed tomorrow.

    EIT recommissioning has been successfully completed. No damage or degradation could be detected. In fact, due to the long and warm bake out, the sensitivity has increased by approximately 60%.

    Recommissioning of the other instruments is progressing smoothly.

18 October 1998

    The spacecraft status is nominal.

    SWAN was successfully switched on at 17:25 UT. All parameters are nominal. Internal heaters of the outer mechanisms have been switched on. SWAN will wait 24 hours until thermal equilibrium is reached before resuming testing of the instrument.

    Recommissioning of the other instruments is progressing smoothly.

17 October 1998

    The spacecraft status is nominal.

    CDS was successfully switched on at 19:55 UT. The Command and Data Handling System and the Experiment Power Supply show nominal telemetry. The Mechanism Subsystem (MCU) was powered on and basic heater circuit checks were successful - CDS heaters remain off for the present with thermal stability being maintained by the s/c heaters until other subsystems have been tested. Both the Slit and Mirror mechanisms have been succesfully tested. Details are available as they happen on the CDS Web page at http://solg2.bnsc.rl.ac.uk

16 October 1998

    The spacecraft status is nominal.

    Momentum management, orbit correction and roll maneuver were successfully completed on 17 October, 03:30 UT.

15 October 1998

    The spacecraft status is nominal.

    UVCS has successfully completed its first O VI mirror scan above the east limb and has successfully taken its Lyman-alpha "re-commissioning first light spectrum". It then has been brought in safe state for the upcoming spacecraft maneuvers. For more details see http://cfa-www.harvard.edu/uvcs/recommission.html.

    The checkout of LASCO/EIT continues to be very successful. Absolutely no problems were encountered during the chechout of the LASCO C2 telescope system. The C2 door was opened and the corona was imaged first in H-alpha, and then with the normal orange filter. The images are excellent. No image degradation is perceived. The C2 and C3 doors then were closed in preparation for the spacecraft maneuvers scheduled for Friday, October 16.

    MDI recommissioning is also progressing successfully. A first Dopplergram was successfully acquired.

14 October 1998

    The spacercraft status is nominal.

    Check out of the LASCO coronagraphs continued. The LASCO C3 door was opened and solar corona images were successfully acquired. The image quality looks fine. An EIT and C3 observing sequence was loaded to run over night.

    MDI continued with software uploads.

    At 00:09 UT UVCS has successfully taken its "recommissioning first light spectrum" in the O VI channel. Looks good! Details are available at http://cfa-www.harvard.edu/uvcs/recommission.html.

13 October 1998

    The spacecraft status is nominal.

    The LION instrument of the COSTEP experiment was successfully switched on at 16:04 UT after MDI had verified that its door is open. All pertaining digital and analog housekeeping channels display characteristics identical to the LION performance prior to the mission interruption, some detector leakage currents show even better (lower) values. Subsequent LION recommissioning and calibration verified that the LION science channels respond to the radiation environment as expected, and no signs of degradation during the unpowered time interval could be detected.

    The LASCO/EIT electronics box was successfully switched on at 17:33 UT and is working well. All four cameras have been loaded and are returning good status. The LASCO C1 Fabry-Perot has also been loaded. Dark images have been taken from all four cameras. The EIT camera has been completly checked. Sun images in all four wavelengths have been taken and they look great! Apparently the 3 months bake-out was very beneficial to the CCD. The latest EIT images are available on the web at: http://umbra.nascom.nasa.gov/eit/eit_full_res.html LASCO recommissioniong will continue over the next several days. Details about the LASCO/EIT recommissioning can be found at: http://lasco-www.nrl.navy.mil/recommiss/recom_sum.html

    MDI turned on the limb tracker and verified open loop functionality. Several additional commissioning test have been performed and an upload of observing programmes has been prepared. MDI took several new images with a warmer filter oven and they look good. For more details see http://quake.stanford.edu/~phil/restart/restart.html

    UVCS also made excellent progress in their recommissioning activities. They successfully recommissioned the occulter mechanisms, the O VI and Ly-alpha grating mechanisms, the O VI and Ly-alpha slit mechanisms, and the roll and pointing mechanisms. The UVCS door was successfully opened and both detectors were powered on. For details see http://cfa-www.harvard.edu/uvcs/recommission.html

    I guess this was a good day.

12 October 1998

    The spacecraft status is nominal.

    Lost guide star and went into RMW mode at 17:26 UT. Back to normal mode at 17:55 UT.

    The long DSN pass on 11 October was lost due to an emergency of TDRS-7. Phase II of the UVCS recommissioning therefore was delayed by a day. UVCS recommissioning continues to proceed smoothly. The door has been partially opened. For further updates see: http://cfa-www.harvard.edu/uvcs/recommission.html

    MDI was successfully switched on at 19:39 UT. The control processor, image processor, camera, and mechanisms (shutter, tuning wheels, calibration wheels) have been successfully switched on. All are functional. The first picture was taken at 22:23 UT. Shutter, camera and image processor worked nominally. The images show faculae and sunspots, and the limb appears sharp, indicating that there is a reasonable light path through the instrument. The images are also well centered. One of the "first light" images is available on the web at: http://mdiwww.nascom.nasa.gov/~zowie/tmp/MDI-recov-2.gif The MDI operators-log is available at http://mdisas.nascom.nasa.gov/cgi-bin/monlog For more information see the SOI home page at: http://soi.stanford.edu/

10 October 1998

    The spacecraft status is nominal.

    UVCS has been successfully switched on at 17:27 UT. Current is nominal and all temperatures are reasonable. NRT commanding from the UVCS EOF workstation works smoothly. A memory checksum test has verified the integrity of the UVCS flight software. At 19:03 UT UVCS was switched to "standby" mode.

    Progress of the UVCS re-commissioning activities can be followed "live" on the UVCS recommissioning page at URL: http://cfa-www.harvard.edu/uvcs/recommission.html

9 October 1998

    The spacecraft status is nominal.

    COSTEP and ERNE (CEPAC) were successfully switched on at 10:33 UT, and both instruments are working fine.

    The turn-on sequence (based solely on TSTOL procedures) went smoothly and resulted in ERNE and COSTEP/EPHIN being in nominal observation mode by 11:30 UT. The LION sensor of COSTEP was not turned on yet, as the MDI door must be open when LION is powered up due to known straylight interference. Temperatures, power consumption, and all instrument housekeeping channels showed nominal values. The common subsystems of ERNE and COSTEP, the low voltage power converter LVPC and the data processing unit CDPU, were also functioning nominally.

    EPHIN recommissioning including calibration, extended calibration, and patch uploads was finished by 14:00 UT. First analysis of science channels shows nominal behaviour, especially all detectors in the particle telescope are alive and display characteristics identical to their performance prior to the loss of the spacecraft. No sign of adverse effects due to thermal stress during the deep freeze have been detected so far.

    The performance of both ERNE sensors, LED and HED, is good. HED has even improved in the sense that one of the detector leakage currents, which was slowly increasing during the flight, has now returned to its initial value. Recommissioning of ERNE, which includes two long calibration sequences, was started at 20:00 UT and is still ongoing.

    GOLF is also back to nominal observing mode. All instrument parameters (except the quarter-wave mechanism) have been returned to previous settings and the instrument sensor is already close to reaching a thermal equilibrium. The daily pulse has been correctly received this night at 00:00. As it requires many hours of NRT commanding to bring the quarter-wave mechanism in the previous position, it was decided to move it to the closer position where the instrument samples the red wing of the sodium lines instead of the blue wing. A decision on the future observing configuration will be taken by the GOLF team after analysis of the data.

8 October 1998

    The spacecraft status is nominal.

    GOLF was successfully switched on at 17:37 UT. All subsytems are functional. A quarter wave mechanism which had moved by approximately 75 deg is being brought back to the nominal position.

    VIRGO announced a preliminary figure for the solar constant: 1365.66 Wm-2

    SUMER is back in safe mode with its door closed. Recommissioning of SUMER will be resumed early November.

7 October 1998

    The spacecraft status is nominal.

    A Fine Pointing Sun Sensor (FPSS) offset of -198 arcsec has been applied to bring the S/C back to the nominal pointing position.

    Further tests on the SUMER-B detector performed last night were successful. The B detector now is working nominally, too.

    VIRGO continues to work fine.

    Tomorrow we will switch on GOLF.

6 October 1998

    The spacecraft status is nominal.

    VIRGO was successfully switched on at 17:55 UT. All instruments (SPM-A, SPM-B, PMO-A, PMO-B, DIARAD, and LOI) have been tested and are working fine. VIRGO is back to nominal.

5 October 1998

    The spacecraft status is nominal.

    SUMER was successfully switched on by the FOT at 18:21 UT. The door was opened through NRT commands at 19:41 UT. Both detectors were switched on. Detector A is working fine, testing of detector B is still ongoing.

2 October 1998

    The spacecraft status is nominal.

    Today the SSU patch was loaded, which was planned for uplink in June to improve the startracker's robustness against single event upsets. The patch did not perform as expected, leading to a transition to RMW mode at 17:46 UT. The spacecraft was back in normal mode at 19:56 UT after re-establishing the previous configuration. The cause for the malfunction of the patch is under investigation.

    Further tests with gyro A did not produce any new results.

    An instrument recommissioning meeting with technical experts and operations personnel from all instrument teams is scheduled for Tuesday, October 6. A draft timeline for the instrument recommissioning is available at: http://sohowww.nascom.nasa.gov/~soc/recommiss.html.

1 October 1998

    Spacecraft status is nominal.

    Problems were encountered during uplink of the star sensor unit (SSU) patch this morning. The problems have been identified and another attempt will be made tomorrow.

    Telemetry subformat #5 has been uplinked and ESR monitoring has been enabled.

30 September 1998

    Spacecraft status remains nominal. Bi-monthly Attitude Control Unit (ACU) memory dumps were performed, an ACU monitoring patch was uplinked, and a test of Gyro C was performed. The Gyro C test indicated that the gyro motor was spinning; however no output from the gyro was received.

    Central Data Management Unit (CDMU) memory dumps continue today. A star sensor unit (SSU) patch, will be performed tomorrow morning.

    Updates of ESR recovery procedures have been defined such that recovery is based on one gyro only. ESR warning flag will be enabled today.

29 September 1998

    Activities yesterday included a change in the guide star. The one being used now is magnitude 5.7, in slot 1. Also, there were group dumps and configuration monitoring. We had some problems while commanding through DSS27, as only 200 Watts uplink power was available. Uploading of COBS and SSU patches was postponed until today, to use DSS24 instead.

    In the morning meeting today there was an overview of the current plans for instrument re-commissioning, with an schedule some 30 working days long.

    Next orbital maneuver is scheduled for October 14.

25 September 1998, evening update

    SOHO is back in normal mode.

    The first activity of the day was a roll trim that was carried out at 13:30 UT. Afterwards, receiver 1 was tested and locked while using the high gain antenna. Receiver 1 is used for ranging in normal mode.

    A station keeping maneuver was completed in two burns along the X axis, direction away from the Sun. The first burn was 2 minutes and the second one 47.5 minutes long. FDF estimates that the resulting delta-vee was within 1% of the desired value of -6.2 m/s.

    After the maneuver was completed, SOHO was commanded to high rate telemetry and then on to normal mode. The spacecraft entered normal mode at 19:52:58 UT.

    A period of two weeks of ranging starts now. Next maneuver, therefore, will be scheduled for some time after October 8.

    Next week will continue the recomissioning activities of spacecraft elements not used during the recovery maneuver. Also, there will be patch uploads and further tests of gyros A and C.

25 September 1998

    The thermal status of both the service and payload modules of the spacecraft is now nominal.

    Receiver 1 was locked twice using the low gain antenna, 18 KW uplink power with a 26m DSN antenna, and wide frequency sweeps.

    A star map was collected and dumped using the star sensor unit 1 (SSU1). FDF reports that SSU1 behaviour is nominal. The absolute roll attitude was determined to be 50.84 degrees.

    SOHO is now in medium rate telemetry, using the high gain antenna (HGA), receiver 2 and transmitter 1. The signal level at DSS24 was -116 dBm, which indicates good HGA pointing.

    Overnight, the tape recorder was set on record during 3 hours, and the data collected were dumped afterwards. The analysis of the dump shows good statistics. The same was done with the solid state recorder, but the analysis of the dump is still under way.

    The procedures for a station keeping maneuver have been prepared and verified.

24 September 1998

    The DHSS problem found the day before yesterday was not a hardware issue, but due to the uplink of Mode 3 commands a little bit too fast. Once the timing of this commands was sligthly changed, the script to transition the spacecraft from ESR mode to RMW (through ISA and FSA modes) was run again yesterday morning without a glitch. The spacecraft is in RMW mode, using a 4.7 magnitude star, since 16:58 UT, yesterday. While on FSA, both SSUs were successfully checked out.

    Other activities accomplished were the uplinking of ACU patches 5 and 6, and the warm start up recovery using the nominal recovery procedure, so DHSS is now back in the nominal branch.

23 September 1998

    Yesterday, there was an attempt to transition the spacecraft from ESR to ISA, then to FSA and on to RMW mode. However, it was not successful for the following two reasons.

    While on ISA mode and preparing the transition to FSA, a Mode 3 command with data words to change the limit for RRAD was sent to the spacecraft. Unfortunately, while building the command, the CMS on the ground shifted the value of this limit 8 bits to the right, dividing it by 256, and making it very close to zero. As soon as this command was uplinked, a roll rate anomaly was triggered on board and the spacecraft transitioned itself to ESR mode again.

    Later in the afternoon, when running the script again, a DHSS error, under investigation, prevented the completion of the activities when it caused a COBS warm start up. This event had no effect on the spacecraft thermal configuration.

    Today, a spacecraft emergency was declared in order to secure DSN coverage until SOHO is, once again, in nominal mode.

    Thrusters 7 and 8 (B side) were successfully tested when SOHO was on ISA mode. This completed the tests of the propulsion subsystem.

22 September 1998

    Yesterday, both Fine Pointing Sun Sensors (FPSS) were checked out and found to be in good health. Also, there were two 0.1 deg/s corrections of the roll rate of the spacecraft.

    The procedures to transition from ESR to ISA and from ISA to RMW are being worked out.

21 September 1998

    On Saturday evening Gyro B was tested and was found to be fine. Gyro C was showing intermittent anomalies in its behaviour. An investigation on this matter is under way. It is worth noting that both gyros with problems, A and C, were warm at the time of the loss of attitude in June while Gyro B was cold.

    The roll rate anomaly detection (RRAD) has been tested also on Saturday and it is working fine. The reaction wheels are being kept at 1500 rpm and the roll rate of the spacecraft never surpassed 0.1 deg/s.

    Plans for testing of thursters 7 and 8 on the B side, and then reconfiguring the spacecraft to the A side are under preparation.

19 September 1998

    Another good day:

      - Successfully spun up all 4 reaction wheels which are now running at 1500 RPM. They behave nominally.
      - Enabled protection and reconfiguration of central onboard computer (CMDU).
      - In case of an anomaly it will now automatically switch over to the redundant side.
      - Successfully performed tape recorder maintenance.
      - Work on gyro B has started (heating).

    Automatic Gain Control (AGC) estimate of the spacecraft roll rate is about -0.1 deg/sec.

18 September 1998

    This was a good day:

      - Gyro C was spun up and is working fine.
      - All thrusters on branch A were successfully tested, including thrusters 7+8.
      - The Solid State Recorder (SSR) was successfully tested.
      - The High Gain Antenna (HGA) pointing mechanism was successfully tested.

17 September 1998

    Spacecraft stability continues to be excellent. The roll rate has been less than 0.2 deg/sec and has now been stabilized to almost zero by thruster firings. An attempt to restart gyro A has been unsuccessful and a reconfiguration to gyro C is planned for tomorrow.

16 September 1998

    The first but important step in the SOHO recovery has been successfully completed today. Sun pointing (without roll control) has been achieved at 18:30 UT. All operations went according to plan. The experiment substitution heaters were switched on 42 minutes after the ESR was triggered.

    During the run up of gyro A an anomaly was detected which is currently under investigation. Meantime the S/C is maintaining the ESR in stable condition with no sign of roll rate being out of limits. The S/C is expected to be stable over night.

    Details of the attitude recovery maneuver run today can be found here.

15 September 1998

    The heating phases of the batteries charging/heating cycles have been reduced by 2 hours. The cycles are now 2 hr charging and 5 hr heating. With these settings the batteries are expected to be fully charged for the attitude recovery maneuver by 01:00 UT tomorrow. Due to this change, both battery temperature and voltage have raised slightly while propulsion subsystem temperatures have decreased a little bit.

    The payload RTU has been switched on briefly today at 16:00 UT to take a reading of the payload module thermistors. The payload temperatures can be found in the table included below.

    After reviewing the simulation run today, it was decided to run the attitude recovery maneuver tomorrow. The solution chosen to recover the attitude of the spacecraft consists of (times are subject to change):

    Time UT Description 1. 06:00 Additional warming up of the propulsion subsystem. 2. 08:00 Intermediate check of telemetry. 3. 10:30 Adjustment of Sun heating mode before maneuvers. 4. 12:00 Thrusters tests. 5. 12:30 Thruster 4B despin calibration. 6. 13:05 Thruster 4B despin 1 (down to 2.4 deg/s). 7. 13:50 Data collection and procedure uploading. 8. 14:10 Data evaluation. 9. FOT load update prior to despin 2. 10. 15:15 Thruster 4B despin 2 (down to 0.9 deg/s). 11. 16:35 Switch on FDE and enable/reset min/max monitoring. 12. 17:09 Data evaluation. 13. Battery management. 14. 17:14 Preparation of ESR. 15. Check spacecraft status before ESR. 16. 18:09 Go for ESR given subsystem by subsystem. 17. 18:10 Thermal control management. 18. Upload timetags to enable standard monitoring functions. T1 ESR trigger. 19. T1+3 ESR without on board roll control. 20. Set up payload configuration (RTU, PPDU on). 21. Trigger COBS reboot and reset thermal control. 22. Thermal adjustment after ESR (including EXP subs heaters). 23. Spin up and selection of gyro for ground roll control.

    Every step has criteria for success and contingency plans defined.

    After the end of these activities, the spacecraft will have on board Sun pointing control. The roll rate will be controlled from ground.

    PLM Temperatures on 9 August, 28 August, 9 September and 15 September 1998 MNEMONIC Limit Temperature [deg C] 8/9/98 8/28/98 9/9/98 15/9/98 CDS QTCI1A 10/30 78.57 48.23 15.14 -4.17 QTCI1B 10/30 78.57 47.60 14.85 -4.47 QTCR1 5/35 1.91 -6.54 -13.98 -18.83 QTCR2 6/30 35.28 13.13 -3.88 -12.70 QTCR3 6/34 41.92 12.56 -9.50 -19.20 CELIAS QTFI1A -15/35 42.78 22.95 6.93 -4.56 QTFI1B -15/35 41.91 22.95 7.32 -3.57 QTFI2A -40/40 35.37 22.95 13.96 1.46 QTFI2B -40/40 34.93 22.95 13.96 1.07 QTFI3A -15/35 41.91 24.12 6.93 -3.07 QTFI3B -15/35 41.91 23.73 6.54 -3.57 QTFR1 5/35 3.35 -4.47 -11.74 -16.63 QTFR2 5/35 3.92 -5.07 -12.38 -18.10 QTFR3 5/35 9.96 -0.68 -9.50 -15.90 QTFR4 7/30 16.74 1.91 -9.50 -17.00 QTFR5 8/36 -5.65 -13.02 -18.83 -21.40* Q11 5/35 10.54 -1.54 -11.10 -17.37 CEPAC QTHI1A 10/25 -21.40* -21.40* -21.40* -21.40* QTHI1B 10/25 -21.40* -21.40* -21.40* -21.40* QTHI2 -13/21 -21.67 -22.22 -24.44 -26.11 QTHI3 0/30 -19.57 -15.58 -16.27 -17.73 QTHR1 5/35 -11.10 -15.58 -20.30 -21.40* QTHR2 8/31 -4.47 -11.74 -18.10 -21.40* QTHR3 -3/34 -21.40* -21.40* -21.40* -21.40* QTHR4 4/33 -11.10 -17.00 -21.40* -21.40* QTHR5 5/35 -7.43 -13.98 -19.57 -21.40* GOLF QTGI1A 5/40 -21.40* -16.27 -18.47 -21.40* QTGI1B 5/40 -21.40* -16.27 -18.10 -21.40* QTGR1 5/35 3.64 -5.06 -12.38 -17.73 QTGR2 10/25 39.39 14.56 -2.40 -11.42 QTGR3 5/35 40.03 12.83 -5.65 -14.94 LASCO/EIT QTLI1A 5/30 33.19 22.95 14.35 8.49 QTLI1B 5/30 33.62 23.72 15.13 9.28 QTLI2A 10/30 -49.20* -49.20* -46.17 -46.17 QTLI2B 10/30 -49.20* -47.68 -44.66 -44.66 QTLI3A -10/20 -49.20* -49.20* -47.69 -46.17 QTLI3B -10/20 -49.20* -49.20* -49.20* -47.69 QTLR1 5/35 -3.58 -11.42 -16.63 -20.67 QTLR2 5/35 -15.26 -20.67 -21.40* -21.40 QTLR3 5/35 4.21 -3.58 -10.46 -15.58 MDI QTMI1A 10/35 -21.40* -18.47 -20.67 -21.03 QTMI1B 10/35 -21.40* -15.58 -17.73 -18.10 QTMI1C 10/41 -21.40* -16.63 -18.83 -19.20 QTMI1D -105/40 -53.33 -43.33 -48.33 -46.67 QTMI2 0/40 -22.15 -23.60 -24.32 -25.05 QTMR1 5/35 -2.40 -9.20 -14.94 -19.57 QTMR2 5/35 -3.58 -8.91 -12.70 -15.90 SUMER QTSI1A 15/27 -21.40* -21.40* -21.40* -21.40* QTSI1B 15/27 -21.40* -21.40* -21.40* -21.40* QTSI2A 5/25 -21.40* -19.93 -21.40* -21.40* QTSI2B 5/25 -21.40* -19.56 -21.03 -21.40* QTSR1 5/35 -0.96 -8.91 -14.94 -19.20 QTSR2 5/34 -16.27 -21.40* -21.40* -21.40* QTSR3 2/27 -13.98 -19.57 -21.40* -21.40* SWAN QTNI1A 5/30 44.45 37.50 30.09 23.04 QTNI1B 5/30 45.71 39.39 31.93 24.90 QTNI2A 5/30 -21.40* -21.40* -21.40* -21.40* QTNI2B 5/30 -21.40* -21.40* -21.40* -21.40* QTNR1 5/35 29.34 20.45 11.11 3.92 QTNR2 5/35 -21.40* -21.40* -21.40* -21.40* QTNR3 6/31 -13.66 -19.57 -21.40* -21.40* UVCS QTUI1A 0/40 -5.65 -6.54 -8.61 -12.06 QTUI1B 0/40 -5.36 -6.25 -8.02 -11.42 QTUR1 5/35 -5.06 -9.82 -15.58 -20.30 QTUR2 5/30 37.13 11.11 -8.32 -17.73 VIRGO QTVR1 3/30 -9.50 -15.58 -21.03 -21.40* QTVR2 5/33 -18.83 -21.40* -21.40* -21.40* ----------- Note: * means: end of sensor measurement range

14 September 1998

    There were no changes in the spacecraft. Propulsion system temperatures remain stable. The heating phases of the charge/heating cycles have been shortened to 7 hours.

    The verification of the procedures for attitude recovery was completed today. A rehearsal of the attitude recovery maneuver is scheduled for tomorrow morning.

12 September 1998

    There were no changes in the spacecraft. Battery charge and temperature, and propulsion system temperatures are stable and being maintained using the 2 hour charging and 8 hour heating cycles started on September 10. This mode will change to two cycles of 2h/5h right before the attitude recovery maneuver to increase the charge of the batteries.

    Receiver 1 was tested twice without accomplishing lock.

    The definition, coding, and verification of attitude recovery and contingency procedures continues.

11 September 1998

    Battery charge and propulsion system temperatures are being maintained. Communication with the spacecraft is carried out using transponder 2 and the +Z low gain antenna. The preparation of the procedures for the attitude recovery maneuver is progressing.

10 September 1998

9 September 1998

    Data was collected from several subsytems in preparation for the attitude recovery maneuver, which is planned for next week. New PLM temperature readings are attached. A test with transponder 1 indicates that the +Z low gain antenna is now also in the field of view of the Earth. Battery charge and propulsion subsystem temperatures are being maintained.

    ---------------------------------------------------------------------- PLM Temperatures on 9 August, 28 August and 9 September 1998 MNEMONIC Limit Temperature [deg C] 8/9/98 8/28/98 9/9/98 CDS QTCI1A 10/30 78.57 48.23 15.14 QTCI1B 10/30 78.57 47.60 14.85 QTCR1 5/35 1.91 -6.54 -13.98 QTCR2 6/30 35.28 13.13 -3.88 QTCR3 6/34 41.92 12.56 -9.50 CELIAS QTFI1A -15/35 42.78 22.95 6.93 QTFI1B -15/35 41.91 22.95 7.32 QTFI2A -40/40 35.37 22.95 13.96 QTFI2B -40/40 34.93 22.95 13.96 QTFI3A -15/35 41.91 24.12 6.93 QTFI3B -15/35 41.91 23.73 6.54 QTFR1 5/35 3.35 -4.47 -11.74 QTFR2 5/35 3.92 -5.07 -12.38 QTFR3 5/35 9.96 -0.68 -9.50 QTFR4 7/30 16.74 1.91 -9.50 QTFR5 8/36 -5.65 -13.02 -18.83 Q11 5/35 10.54 -1.54 -11.10 CEPAC QTHI1A 10/25 -21.40* -21.40* -21.40* QTHI1B 10/25 -21.40* -21.40* -21.40* QTHI2 -13/21 -21.67 -22.22 -24.44 QTHI3 0/30 -19.57 -15.58 -16.27 QTHR1 5/35 -11.10 -15.58 -20.30 QTHR2 8/31 -4.47 -11.74 -18.10 QTHR3 -3/34 -21.40* -21.40* -21.40* QTHR4 4/33 -11.10 -17.00 -21.40* QTHR5 5/35 -7.43 -13.98 -19.57 GOLF QTGI1A 5/40 -21.40* -16.27 -18.47 QTGI1B 5/40 -21.40* -16.27 -18.10 QTGR1 5/35 3.64 -5.06 -12.38 QTGR2 10/25 39.39 14.56 -2.40 QTGR3 5/35 40.03 12.83 -5.65 LASCO/EIT QTLI1A 5/30 33.19 22.95 14.35 QTLI1B 5/30 33.62 23.72 15.13 QTLI2A 10/30 -49.20* -49.20* -46.17 QTLI2B 10/30 -49.20* -47.68 -44.66 QTLI3A -10/20 -49.20* -49.20* -47.69 QTLI3B -10/20 -49.20* -49.20* -49.20* QTLR1 5/35 -3.58 -11.42 -16.63 QTLR2 5/35 -15.26 -20.67 -21.40* QTLR3 5/35 4.21 -3.58 -10.46 MDI QTMI1A 10/35 -21.40* -18.47 -20.67 QTMI1B 10/35 -21.40* -15.58 -17.73 QTMI1C 10/41 -21.40* -16.63 -18.83 QTMI1D -105/40 -53.33 -43.33 -48.33 QTMI2 0/40 -22.15 -23.60 -24.32 QTMR1 5/35 -2.40 -9.20 -14.94 QTMR2 5/35 -3.58 -8.91 -12.70 SUMER QTSI1A 15/27 -21.40* -21.40* -21.40* QTSI1B 15/27 -21.40* -21.40* -21.40* QTSI2A 5/25 -21.40* -19.93 -21.40* QTSI2B 5/25 -21.40* -19.56 -21.03 QTSR1 5/35 -0.96 -8.91 -14.94 QTSR2 5/34 -16.27 -21.40* -21.40* QTSR3 2/27 -13.98 -19.57 -21.40* SWAN QTNI1A 5/30 44.45 37.50 30.09 QTNI1B 5/30 45.71 39.39 31.93 QTNI2A 5/30 -21.40* -21.40* -21.40* QTNI2B 5/30 -21.40* -21.40* -21.40* QTNR1 5/35 29.34 20.45 11.11 QTNR2 5/35 -21.40* -21.40* -21.40* QTNR3 6/31 -13.66 -19.57 -21.40* UVCS QTUI1A 0/40 -5.65 -6.54 -8.61 QTUI1B 0/40 -5.36 -6.25 -8.02 QTUR1 5/35 -5.06 -9.82 -15.58 QTUR2 5/30 37.13 11.11 -8.32 VIRGO QTVR1 3/30 -9.50 -15.58 -21.03 QTVR2 5/33 -18.83 -21.40* -21.40* ----------- Note: * means: end of sensor measurement range

8 September 1998

    The batteries are fully charged. The propulsion subsystem temperatures are slowly improving.

    Three more experts from Europe have arrived at GSFC for the preparation of the attitude recovery. A review of the recovery scenario for final approval is planned for Thursday, 10 September 1998.

7 September 1998

    Battery recharging continues. Propulsion subsystem temperatures are being maintained.

    Two alternative solutions which use fewer thrusters for attitude recovery are being studied:

    Solution A consists of:

      - Thruster test
      - Despin of S/C to low rate around Z-axis
      - Triggering of ESR without gyro roll control

    Solution B consists of:

      - Thruster test
      - Despin of S/C to low rate around Z-axis
      - Spin-up around X-axis while X-axis stays oriented towards Sun
      - Stabilization of attitude by spinning up reaction wheels

4 September 1998

    We are continuing to slowly recharge the batteries whilst trying to maintain the propulsion subsystem temperatures, which is proving to be a delicate balance.

    Alternative scenarios for the attitude recovery with reduced requirements on the propulsion and power subsystems are currently being studied.

3 September 1998

    We are entering a cyclic battery charging phase while trying to maintain propulsion subsystem temperatures.

2 September 1998

    Pipe heating continues. All the pipe heaters are currently switched on in sun-heating mode. Three of the Dual Thruster Modules (DTMs) appear to be thawed. Pressures in the tank and two propulsion branches are as predicted and steady.

    A test was performed on receiver #1 on the +Z low gain antenna. Lock was not achieved.

    A four step approach has been adopted for the attitude recovery:

    1. Gyro test 2. Thruster test 3. Despin of the S/C to low rate 4. Triggering of ESR

    Details are being discussed by AOCS experts at GSFC, MMS-F Toulouse and MMS-UK Bristol. The schedule for these activities depends on the completion of thawing of the propulsion system and the time needed to fully recharge the batteries.

1 September 1998

    Pipe heating is continuing. Three of four pipe sections of the propulsion subsystem are presently being heated while the tank temperature is being maintained at 10.5 deg C. Data from the propulsion instrumentation indicates nominal tank pressure. Propulsion branch pressures are as expected. Latch valve B has been closed to isolate pipes from the tank.

    A test with transmitter 1 on the +Z low gain antenna has been successful. Carrier signal was received for approximately 1 min 45 sec with expected fluctuations due to spacecraft spin.

31 August 1998

    Heating of the first of four pipe sections continues. Pipe temperature and pressure are monitored in one hour intervals. The heating rate is being adjusted accordingly. Tank temperature is being maintained during pipe heating.

30 August 1998

    Heating of the first of four pipe sections has commenced at 12:30 UT this morning after recharging the batteries for 37 hours. Tank temperature is 9.7 deg C. Shift working has been resumed to enable 24 hour monitoring of the pipe temperatures.

28 August 1998

    Continued tank heating. Tank temperature reached 9 deg C at 19:02 UT. Will stop tank heating at 23 UT to recharge batteries for 36 hours. Will start pipe heating Sunday morning.

    The PRTU was briefly switched on at 15:05 UT to obtain another reading of the instrument temperatures (see attachment).

    ---------------------------------------------------------------------- PLM Temperatures on 9 August 1998 and 28 August 1998 MNEMONIC Limit Temperature [deg C] 8/9/98 8/28/98 CDS QTCI1A 10/30 78.57 48.23 QTCI1B 10/30 78.57 47.60 QTCR1 5/35 1.91 -6.54 QTCR2 6/30 35.28 13.13 QTCR3 6/34 41.92 12.56 CELIAS QTFI1A -15/35 42.78 22.95 QTFI1B -15/35 41.91 22.95 QTFI2A -40/40 35.37 22.95 QTFI2B -40/40 34.93 22.95 QTFI3A -15/35 41.91 24.12 QTFI3B -15/35 41.91 23.73 QTFR1 5/35 3.35 -4.47 QTFR2 5/35 3.92 -5.07 QTFR3 5/35 9.96 -0.68 QTFR4 7/30 16.74 1.91 QTFR5 8/36 -5.65 -13.02 Q11 5/35 10.54 -1.54 CEPAC QTHI1A 10/25 -21.40* -21.40* QTHI1B 10/25 -21.40* -21.40* QTHI2 -13/21 -21.67 -22.22 QTHI3 0/30 -19.57 -15.58 QTHR1 5/35 -11.10 -15.58 QTHR2 8/31 -4.47 -11.74 QTHR3 -3/34 -21.40* -21.40* QTHR4 4/33 -11.10 -17.00 QTHR5 5/35 -7.43 -13.98 GOLF QTGI1A 5/40 -21.40* -16.27 QTGI1B 5/40 -21.40* -16.27 QTGR1 5/35 3.64 -5.06 QTGR2 10/25 39.39 14.56 QTGR3 5/35 40.03 12.83 LASCO/EIT QTLI1A 5/30 33.19 22.95 QTLI1B 5/30 33.62 23.72 QTLI2A 10/30 -49.20* -49.20* QTLI2B 10/30 -49.20* -47.68 QTLI3A -10/20 -49.20* -49.20* QTLI3B -10/20 -49.20* -49.20* QTLR1 5/35 -3.58 -11.42 QTLR2 5/35 -15.26 -20.67 QTLR3 5/35 4.21 -3.58 MDI QTMI1A 10/35 -21.40* -18.47 QTMI1B 10/35 -21.40* -15.58 QTMI1C 10/41 -21.40* -16.63 QTMI1D -105/40 -53.33 -43.33 QTMI2 0/40 -22.15 -23.60 QTMR1 5/35 -2.40 -9.20 QTMR2 5/35 -3.58 -8.91 SUMER QTSI1A 15/27 -21.40* -21.40* QTSI1B 15/27 -21.40* -21.40* QTSI2A 5/25 -21.40* -19.93 QTSI2B 5/25 -21.40* -19.56 QTSR1 5/35 -0.96 -8.91 QTSR2 5/34 -16.27 -21.40* QTSR3 2/27 -13.98 -19.57 SWAN QTNI1A 5/30 44.45 37.50 QTNI1B 5/30 45.71 39.39 QTNI2A 5/30 -21.40* -21.40* QTNI2B 5/30 -21.40* -21.40* QTNR1 5/35 29.34 20.45 QTNR2 5/35 -21.40* -21.40* QTNR3 6/31 -13.66 -19.57 UVCS QTUI1A 0/40 -5.65 -6.54 QTUI1B 0/40 -5.36 -6.25 QTUR1 5/35 -5.06 -9.82 QTUR2 5/30 37.13 11.11 VIRGO QTVR1 3/30 -9.50 -15.58 QTVR2 5/33 -18.83 -21.40* ----------- Note: * means: end of sensor measurement range

27 August 1998

    Resumed tank heating at 15:03 UT after charging the batteries for 29 hours. Tank temperature at 19:05 UT: 7 deg C. Will continue tank heating until it reaches 10 deg C (expected for Saturday morning). Presently we plan to start pipe heating Sunday morning after topping up the batteries.

26 August 1998

    The tank temperature continued to increase over night:
    02:00 UT: 3.0 deg C
    06:00 UI: 3.9 deg C
    10:00 UT: 5.3 deg C
    Switched of tank heating at 10:10 UT to recharge the batteries for 24 hours.

25 August 1998

    Today appears to be the first sign of the completion of the hydrazine thawing of the tank: The tank temperature increased from 1.3 deg C to 1.9 and 2.3 deg C at 10:00 and 18:00 UT, respectively. Tank heating will continue until tomorrow morning at 10:00 UT when we will start recharging the batteries.

    ESA and MMS AOCS experts arrived today at GSFC for the preparation of the attitude recovery.

    Colleagues from the simulation section of ESA/ESTEC (TOS-EMM) have prepared two animations showing SOHO's estimated attitude on 23 August (45 deg) and 7 September (60 deg): http://sohowww.nascom.nasa.gov/operations/ESR7/attitude.html

24 August 1998

    Tank heating continues.

23 August 1998

    Stopped recharging the batteries and resumed tank heating at 02:00 UT.

21 August 1998

    Stopped tank heating at 20:15 UT to recharge batteries for 30 hours. Will resume tank heating tomorrow night. Once the propellant in the tank has been brought to 10 deg C, we will start heating the pipes. After pipe heating, which will take about 2 days, there will be a gyro test, HPA test (in high power mode) and thruster test. The schedule for these activities is still under discussion.

    Prelmininary analysis of the ranging data indicates that we are very close to the predicted orbit before the momentum management on 24 June, i.e. the delta V disturbances experienced during ESR 5, 6 and 7 were very small. To give you an indication of how good the orbit still is: For an orbit correction on 15 September we would need only a few kg of hydrazine.

20 August 1998

    Tank heating continues.

19 August 1998

    Thawing of the tank is progressing. A new procedure to heat only in sunlight without discharging the batteries has been successfully tested. This procedure will be applied later to maintain thermal control of the propulsion subsystem while charging the batteries.

18 August 1998

    Stopped recharging the batteries at 20:09 UT and resumed heating of the tank at 20:15 UT. Took 20 min of SAS-A data starting at 20:25 UT. Battery charge state and temperatures will be checked in 6 hour intervals.

17 August 1998

    Thawing of the tank proceeded as planned. Heating was stopped at 19:05 UT to recharge the batteries.

16 August 1998

    Thawing of the tank is proceeding as planned. A short switch-on test of the FDE (Failure Detection Electronics) has been successful.

15 August 1998

    Thawing of the tank is proceeding as planned.

14 August 1998

    Charging of the batteries continued until 20:05 UT. They reached the expected value of 46 V (approx. 65% of max capacity). At 20:17 UT the heaters of the tank were switched on again to continue thawing of the hydrazine. Temperatures and battery voltage are monitored in 4 hour intervals.

13 August 1998

    The SAS data received on 11 August have been analyzed. It was confirmed that SOHO is spinning around an axis close to the Z axis at a rate of -6.9 deg/sec (rotation period 52.6 sec). The rotation vector is 1.8 deg from the Z axis towards the X axis and 7.6 deg from the Z axis towards the -Y axis. The angle between the rotation axis and the Sun is 36.7 deg.

    The X-band ranging on 11 August was not successful. The cause for the failure is still under investigation.

    Today we started ranging with SOHO. We had a total of four session where we acquired ranging data, each lasting a couple of minutes.

    Thawing of the tank has started yesterday evening at 22:39 UT. It was interrupted today at 20:03 UT in order to recharge the batteries. Analysis of the power data revealed a slightly negative power balance, i.e. we draw power from the batteries when the heaters are on. We therefore have to recharge the batteries from time to time. Batteries charged for 20 hours should have enough capacity to allow us to turn the heaters on for about 3 days. Presently we estimate that about two cycles of charging/heating will be necessary to thaw the hydrazine in the tank.

12 August 1998

    Power, thermal, and propulsion experts from ESA/ESTEC, MMS-F and MMS-UK have arrived this morning.

    A detailed power budget has been established and a decision has been made to start thawing of the propellant in the tank this evening. The batteries' charge state and tank temperature will be carefully monitored to confirm power budget and thermal estimates. This monitoring required us to declare a spacecraft emergency in order to have continuous DSN contact.

    The SAS data acquired yesterday are still being analyzed at ESTEC, MMS-F and GSFC. Also no definitive results yet from yesterday's X-band ranging measurement.

11 August 1998

    The S-band radar measurements planned for today could not be performed due to techical problems with the Arecibo transmitter. In the afternoon, DSS14 performed X-band ranging measurements for about an hour. At the time of writing of this report it was not clear whether we succeeded to obtain valid data.

    At 21 UT we successfully switched on Attitude Control Unit A (ACU-A) and Control Actuation Electronics A (CAE-A) which allowed us to get attitude data from the Sun Acquisition Sensors (SAS). 5 min of SAS data were acquired. Experts will analyze the data tonight to determine the exact attitude of the spacecraft. The telemetry session is still going on at this time (now for almost 1 1/2 hours). The FOT is monitoring the power status of the batteries, which seem to hold the power well.

    If the power balance proofs to be adequate, we will start to thaw out the propulsion subsytem tomorrow.

10 August 1998

    On Sunday afternoon we had two more successful telemetry sessions, both with PRTU ON. To preserve resources, the first session was terminated through ground commands after 4 min 15 sec, in the second session telemetry was switched off after 5 min 1 sec. Experts at GSFC, ESA/ESTEC and MMS are analyzing the data. As expected, the spacecraft in general is cold. Of course, some units are warmer than normal. The tank is around 1 deg C, so the hydrazine probably is partially frozen.

    Attached you can find temperatures for the PLM units. Those annotated with a star have no meaning (out of sensor measurement range).

    Presently we are charging both batteries. As a safety precaution, charging of the batteries has been limited to 45 V.

    Propulsion, power, and AOCS experts are expected to arrive at GSFC later this week.

    PLM Temperatures on 9 August 1998 MNEMONIC Limit Temperature [deg C] CDS QTCI1A 10/30 78.57 QTCI1B 10/30 78.57 QTCR1 5/35 1.91 QTCR2 6/30 35.28 QTCR3 6/34 41.92 CELIAS QTFI1A -15/35 42.78 QTFI1B -15/35 41.91 QTFI2A -40/40 35.37 QTFI2B -40/40 34.93 QTFI3A -15/35 41.91 QTFI3B -15/35 41.91 QTFR1 5/35 3.35 QTFR2 5/35 3.92 QTFR3 5/35 9.96 QTFR4 7/30 16.74 QTFR5 8/36 -5.65 Q11 5/35 10.54 CEPAC QTHI1A 10/25 -21.40* QTHI1B 10/25 -21.40* QTHI2 -13/21 -21.67 QTHI3 0/30 -19.57 QTHR1 5/35 -11.10 QTHR2 8/31 -4.47 QTHR3 -3/34 -21.40* QTHR4 4/33 -11.10 QTHR5 5/35 -7.43 GOLF QTGI1A 5/40 -21.40* QTGI1B 5/40 -21.40* QTGR1 5/35 3.64 QTGR2 10/25 39.39 QTGR3 5/35 40.03 LASCO/EIT QTLI1A 5/30 33.19 QTLI1B 5/30 33.62 QTLI2A 10/30 -49.20* QTLI2B 10/30 -49.20* QTLI3A -10/20 -49.20* QTLI3B -10/20 -49.20* QTLR1 5/35 -3.58 QTLR2 5/35 -15.26 QTLR3 5/35 4.21 MDI QTMI1A 10/35 -21.40* QTMI1B 10/35 -21.40* QTMI1C 10/41 -21.40* QTMI1D -105/40 -53.33 QTMI2 0/40 -22.15 QTMR1 5/35 -2.40 QTMR2 5/35 -3.58 SUMER QTSI1A 15/27 -21.40* QTSI1B 15/27 -21.40* QTSI2A 5/25 -21.40* QTSI2B 5/25 -21.40* QTSR1 5/35 -0.96 QTSR2 5/34 -16.27 QTSR3 2/27 -13.98 SWAN QTNI1A 5/30 44.45 QTNI1B 5/30 45.71 QTNI2A 5/30 -21.40* QTNI2B 5/30 -21.40* QTNR1 5/35 29.34 QTNR2 5/35 -21.40* QTNR3 6/31 -13.66 UVCS QTUI1A 0/40 -5.65 QTUI1B 0/40 -5.36 QTUR1 5/35 -5.06 QTUR2 5/30 37.13 VIRGO QTVR1 3/30 -9.50 QTVR2 5/33 -18.83 ----------- Note: * means: end of sensor measurement range

9 August 1998

    SOHO telemetry has been restored on Saturday, 8 August 1998 at 23:15 UT. After recharging battery #2 for 10 hours, telemetry was successfully switched on and seven frames of VC0 have been received at GSFC from Goldstone (D24). Telemetry has been switched off after 1 minute to preserve resources. The temperature of the service module is as expected very low. No P/L temperatures are available yet. We are discussing the possibility to try and acquire telemetry today with PRTU ON to get a snapshot of the P/L temperatures. The parts of the power, data handling and telecommunication subsystems used so far behaved nominally. Battery #2 is partially charged (42 V; nominal 46.8 V). Experts in Europe have been contacted for further advice on the management of the batteries.

7 August 1998

    The Goldstone station has been reconfigured so that partial frames could be recorded and sent out to Goddard. During a 1 hour test late this afternoon we got carrier lock, subcarrier lock and symbol lock for 5 sec. The validity of the data has to be confirmed after processing. The test was limited in order to limit the on/off stress of the units involved.

    A new procedure to charge the batteries will be run over night.

    We were able to get time at Arecibo and DSS14 next Tuesday, August 11. Arecibo will support SOHO radar observations from 1300 to 1600 UT. DSS14 is scheduled for SOHO from 1145 to 2115 UT. The plan is to do S-band bistatic measurements with Arecibo in the slot 1145-1600, and X-band radar ranging from 1600 to 2115 UT. Arecibo will let us know Monday whether the S-band transmitter will be operational. If Arecibo is not operational, the time will be used for SOHO TLM downlink acquisition efforts.

6 August 1998

    With D24 send/receive and D27 receiving we got many more signal spikes, some of which lasted up to 60 sec. Data modulation on the telemetry subcarrier was confirmed. However, we still have not been able to lock on the subcarrier and demodulate the signal. Efforts are being made to validate the ground system.

    A test was performed using Transmitter 1 (which is connected to the +z low gain antenna). No signal was received. (Commanding through Receiver 2)

    A test with carrier ON only produced spikes of the expected signal strength (-135 dbm).

    A test with the previously unsuccessful switch-on procedure now was successful.

    Tom van Overbeek and colleagues from the simulation section of ESA/ESTEC (TOS-EMM) have produced an MPEG animation of the time period 100-260 sec of the MMS-UK simulation of ESR7 which matches the available telemetry the best way. You can find the MPEG and some other information about the last minutes before the loss of telemetry on June 25 at: http://sohowww.nascom.nasa.gov/operations/ESR7/ or, for our European colleagues, at our mirror site at the Solar Sytem Divsion at ESTEC (SCI-SO): http://sohowww.estec.esa.nl/operations/ESR7/

5 August 1998

Today we succeeded to switch on telemetry modulation for periods lasting 2 to 14 sec. Occasionally, the ground station receiver achieved lock, but the duration was too short to decode the telemetry data. Discussions are ongoing with JPL to find ways to decode shorter bursts of telemetry.

Unfortunately, the radar measurements planned for today could no be performed because a high voltage cable to the Arecibo transmitter in the carriage house has broken. DSS14, which was scheduled to receive Arecibo's signals in bistatic mode, could not be used for monostatic radar measurements because they had no clearance for radar uplink.

4 August 1998

The search finally was successful!

Last evening, August 3, 22:51 UT, after yet another modification of the uplink command sequence, carrier spikes lasting between 2 and 10 sec were received both by DSS42 in Canberra and the ESA station in Perth. >From the repetition rate of the spikes one can deduce a rotation period of about 52 sec. This is in excellent agreement with results reported today by Gregory Black from Cornell, who has done more analysis of the bistatic radar measurements from 23 July. From a Fourier analysis of these data he has extracted the rotation frequency (53sec) and associated harmonics. Two new plots including descriptions are available on his web page at: http://astrosun.tn.cornell.edu/radargroup/SOHO/

The next steps are to recharge the batteries and try to switch on telemetry modulation.

DSN coverage:

Date DOY Time Station 4/8 216 0835 - 1245 DSS61 Madrid 4/8 216 1225 - 0030 DSS24 Goldstone

3 August 1998

The search continues.

DSN coverage:

Date DOY Time Station 1/8 213 0420 - 0920 DSS61 Madrid 1/8 213 1215 - 2300 DSS13 Goldstone (WBSA) 1/8 213 1215 - 0015 DSS24 Golstone 2/8 214 0030 - 0430 DSS42 Canberra 2/8 214 0425 - 0715 DSS61 Madrid 2/8 214 1220 - 2045 DSS24 Goldstone 2/8 214 2115 - 0430/215 DSS34 Canberra 3/8 215 0430 - 1020 DSS61 Madrid 3/8 215 1225 - 1735 DSS24 Goldstone 3/8 215 2245 - 0500/216 DSS42 Canberra

31 July 1998

The search continues.

The Arecibo group is reworking their configuration to attempt a monostatic CW collection next Wednesday, August 5. We are trying to schedule DSS14 at Goldstone to have a backup in case the Arecibo transmit/receive system has problems. In this context one should note that the use of Arecibo to both transmit and receive the echo is complicated by the very short, 10 sec, round trip light time. The Arecibo radar was designed to look at objects at much larger distances so it uses a CW transmitter which takes several seconds to change from transmitting to receiving or vise versa and which involves rotation of a device weighing several tousand pounds. However, if the attempt is successful, it will provide the best data to extract information on the attitude state of SOHO.

DSN coverage:

Date DOY Time Station 31/7 212 0120 - 0430 DSS42 Canberra 31/7 212 0420 - 0545 DSS61 Madrid 31/7 212 1215 - 2125 DSS24 Goldstone 31/7 212 2115 - 0425/213 DSS34 Canberra

30 July 1998

The search continues.

DSN coverage:

Date DOY Time Station 30/7 211 0415 - 0710 DSS61 Madrid 30/7 211 1210 - 0030/212 DSS24 Goldstone

29 July 1998

The search continues.

The 70 m dish of DSS14 at Goldstone was used for monostatic radar measurements in X-band (8.8 GHz) with an uplink power of 150 kW. Again, a good signal was received at the predicted location. Unfortunately, no ranging measurements could be performed because the ephemeris product needed for the pulse mode operation contained an unexpected character, causing the software to stall. As I understood the colleagues from JPL, this format has been used now for 20 years, and yet they discovered another failure mode! We hope to get 70 m time next week to repeat the ranging measurements. Because of this software problem, only CW (continuous wave) measurements could be performed. Preliminary analysis of these data again are consistent with a spin rate of 1 RPM.

The analysis of last week's bistatic radar measurements Arecibo-Goldstone is still ongoing. There are several questions which have not yet been conclusively answered. Among them is that of the significance of the spurious slow drift which I've mentioned in my status report of 23 July. This drift seems symmetric with respect to transit at Arecibo. Some people therefore are suspicious whether this is a real effect or related to an instrumental effect. Further, re-calibration of the data by Greg Black from the NAIC group indicates rather large radar cross sections on the order of 15 to 20 m^2, as compared to previous estimates which were on the order of 4 m^2. These somewhat conflicting results may give you a good indication where we stand with the analysis of last weeks measurements.

Greg Black from NAIC was kind enough to put some of his spectra on the web: http://astrosun.tn.cornell.edu/radargroup/SOHO/soho.html

Please keep in mind that these are still PRELIMINARY RESULTS!

DSN coverage:

Date DOY Time Station 29/7 210 0415 - 0730 DSS61 Madrid 29/7 210 1210 - 1540 DSS24 Goldstone 29/7 210 1410 - 1800 DSS14 Goldstone 29/7 210 2100 - 0410 DSS42 Canberra

28 July 1998

The search continues.

Nothing definitive yet from the analysis of the radar data. We hope to hear more tomorrow evening in a telecon with JPL after the monostatic run with DSS14 in X-band.

DSN coverage:

Date DOY Time Station 28/7 209 0030 - 0345 DSS42 Canberra 28/7 209 0415 - 0730 DSS61 Madrid 28/7 209 1205 - 2355 DSS24 Goldstone 28/7 209 2335 - 0320/210 DSS42 Canberra

27 July 1998

The search continues.

The radar measurements from last week are still being analyzed. Monostatic radar observations in X-band (8.8 GHz) using DSS14 (70 mantenna at Goldstone) are scheduled for Wednesday, July 29, 1240 - 1830 UT.

DSN coverage:

Date DOY Time Station 25/7 206 0405 - 0655 DSS61 Madrid 25/7 206 1200 - 0030/207 DSS24 Goldstone 26/7 207 0010 - 0240 DSS42 Canberra 26/7 207 0405 - 0830 DSS61 Madrid 26/7 207 1200 - 0015/208 DSS24 Goldstone 27/7 208 0410 - 0845 DSS61 Madrid 27/7 208 1205 - 0040/209 DSS24 Goldstone

24 July 1998

The search continues.

Radar experts at NAIC (National Astronomy and Ionosphere Center, Cornell University), JPL and the Lincoln Laboratories are working on the data obtained yesterday. Some sample spectra made at the Goldstone antenna site just after the data were taken are available on the web at: http://astrosun.tn.cornell.edu/radargroup/SOHO/ (They are gif screen captures; one is an animation).

DSN coverage:

Date DOY Time Station 24/7 205 0120 - 0430 DSS42 Canberra 24/7 205 0315 - 0800 DSS61 Madrid 24/7 205 1200 - 2115 DSS24 Goldstone 24/7 205 2100 - 0430 DSS42 Canberra

23 July 1998

The bistatic radar operation Arecibo-Goldstone, which was performed today from 1300 to 1545 UT, seems to have been successful. DSS14 at Goldstone was able to receive strong echos from SOHO at its predicted location. The signal width was 1-2 Hz, which is compatible with a spin rate of about 1 RPM (certainly not much faster). The center frequency drifted slowly by a few Hz. The period of this slow movement must be longer than 2 hours. It may be indicative of a non-principal axis rotation.

All this results from a very preliminary analysis of data gathered just 6 hours ago. Two teams at JPL and Arecibo are working in parallel on a detailed analysis of the data, hopefully leading to a better understanding of SOHO's attitude. We hope to have more to report by Monday.

In the meantime, we are trying to schedule a monostatic DSN run in X-band. With half an hour integration time it should be possible to determine the position to +- 20 m. Periodic measurements of this kind should help keep SOHO's ephemeris up-to-date.

DSN coverage:

Date DOY Time Station 23/7 204 0415 - 0730 DSS61 Madrid 23/7 204 1155 - 2340 DSS24 Goldstone (34 m) 23/7 204 1300 - 1545 DSS14 Goldstone (70 m; bistatic radar with Arecibo)

22 July 1998

The search continues.

Tomorrow, 22 July, 1300-1600 UT, there will be a bistatic radar operation Arecibo - Goldstone. The aims of this operation are - to confirm the location of SOHO - provide updates to orbital vectors - possibly provide information on the spin rate and orientation of the spin axis

DSN coverage:

Date DOY Time Station 22/7 203 0415 - 0750 DSS66 Madrid 22/7 203 1155 - 1445 DSS24 Goldstone 22/7 203 2040 - 0430/204 DSS42 Canberra

21 July 1998

The search continues.

Studies indicate that it may not be unfeasible to locate SOHO using Arecibo radar. On Thursday we will give it a first try using bistatic radar Arecibo - Goldstone. Arecibo has a 1 MW transmitter operating at 2.380 GHz. Goldstone will try to receive the reflected signal with a 70 m dish (DSS14). The team at Arecibo is led by Donald Campbell from Cornell. For more information on Arecibo, see http://www.naic.edu/open.htm

DSN coverage:

Date DOY Time Station 21/7 202 0355 - 0545 DSS61 Madrid 21/7 202 1155 - 2325 DSS24 Goldstone (+DSS13) 21/7 202 2305 - 0405/203 DSS34 Canberra

20 July 1998

The search continues.

A digital Wide Band Spectrum Analyser (WBSA) which was previously used for the SETI project has been installed at Goldstone. In order to be able to eliminate instrumental effects, it has been set up with two simultaneous feeds from DSS24 and DSS13.

DSN coverage:

Date DOY Time Station 18/7 199 1155 - 2345 DSS24 Goldstone 19/7 200 0140 - 0425 DSS34 Canberra 19/7 200 1155 - 2145 DSS24 Goldstone 19/7 200 2130 - 0345/201 DSS34 Canberra 20/7 201 0355 - 0800 DSS61 Madrid 20/7 201 1155 - 0020 DSS24 Goldstone (+ DSS13)

17 July 1998

The search continues.

Colleagues operating the Stanford/SRI 150-feet dish joined in the search for a signal from SOHO. Feasibility studies of locating SOHO using Arecibo and DSN radar are ongoing.

DSN coverage:

Date DOY Time Station 17/7 198 0445 - 0810 DSS61 Madrid 17/7 198 1245 - 2220 DSS24 Goldstone 17/7 198 2210 - 0245/199 DSS42 Canberra

16 July 1998

The search continues.

ESA and NASA have issued a press release "Efforts to recover SOHO spacecraft continue as investigation board focuses on most likely causes". The ESA-NASA Mission Interruption Review Board has issued a preliminary status and background report. Both the press release and the preliminary report are available on the web at http://sohowww.nascom.nasa.gov/whatsnew/

DSN coverage:

Date DOY Time Station 16/7 197 0540 - 0830 DSS61 Madrid 16/7 197 1155 - 2345 DSS24 Goldstone 16/7 197 2320 - 0245/198 DSS42 Canberra

15 July 1998

The search continues.

A system is being installed which allows to view the spectrum of the downlink in real time at the EOF.

DSN coverage:

Date DOY Time Station 15/7 196 0010 - 0420 DSS42 Canberra 15/7 196 0755 - 1135 DSS61 Madrid 15/7 196 1155 - 1445 DSS24 Goldstone 15/7 196 2100 - 0245/197 DSS42 Canberra

14 July 1998

The search continues.

This morning RF experts from ESA/ESOC and JPL have arrived to help optimize the search strategy and the set-up of special equipment at the ground stations.

DSN coverage:

Date DOY Time Station 14/7 195 0015 - 0245 DSS42 Canberra 14/7 195 0830 - 1215 DSS61 Madrid 14/7 195 1155 - 2250 DSS24 Goldstone

13 July 1998

DSN stations are continuing to transmit telecommands to RCVR2 at 2067.221 MHz -> 2067.271 MHz in 1.0 kHz steps. Search for downlink signal in frequency interval 2244.945 +- 0.065 MHz.

DSN coverage:

Date DOY Time Station 11/7 192 0710 - 0945 DSS61 Madrid 11/7 192 1150 - 0000 DSS24 Goldstone 11/7 192 2340 - 0210/193 DSS42 Canberra 12/7 193 0715 - 1000 DSS61 Madrid 12/7 193 1255 - 0000 DSS24 Goldstone 13/7 194 0500 - 0805 DSS61 Madrid 13/7 194 1155 - 0000 DSS24 Goldstone

The search for a downlink signal at two ESA stations in Perth, Australia and Villafranca, Spain is also continuing, so far without success.

 
 

• European Site • US Site

 

Feedback & Comments: SOHO Webmaster

Last modification:

SOHO
and
SOHO
SOHO is a project of international cooperation between